# Non-isogenous superelliptic jacobians II

### v1

@@ 2.1. Galois properties
In this subsection we recall an explicit description of a certain important Galois submodule of $J^{(f,\ell)}[\ell]$ [10, 11] for arbitrary separable $f(x)$ and odd $n$ under an additional assumption that $\ell$ does not divide $n$.
@@ 3. Isogenous superelliptic jacobians
Let $\ell$ be a prime, $F$ a field of prime characteristic $\ell$, and $V$ a finite-dimensional vector space over $F$. Let $\mathcal{N}\subset\mathrm{End}_{F}(V)$ be a linear nilpotent Lie subalgebra of $\mathrm{End}_{F}(V)$, and
@@
    $K(X[\ell])/K(\mathfrak{R}_{f})$ is a finite Galois $\ell$-extension, i.e., either $K(X[\ell])=K(\mathfrak{R}_{f})$ or the Galois group $\mathrm{Gal}(K(X[\ell])/K(\mathfrak{R}_{f}))$ is a finite $\ell$-group.
@@
$\frac{2\mathrm{dim}(J^{(f,\ell)})}{[{\mathbb{Q}}(\zeta_{\ell}):{\mathbb{Q}}]}=n-1.$
@@
is a ring embedding that makes $X[\ell]$ a free ${\mathbb{Z}}[\delta_{\ell}]/\ell$-module. This implies that
@@
Now (i) follows from Lemma 3.3 applied to
@@
Proof of Theorem 2.2 (modulo Lemma 3.6).
@@
Both endomorphism algebras $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(Y)$ contain an invertible element of multiplicative order $r$. This proves Theorem 2.2. ∎
@@
$G_{\ell}:=\tilde{G}_{\ell,X}=\mathrm{Gal}(K(X[\ell])/K),\ G=\mathrm{Gal}(K(X[1-\delta_{\ell}])/K)=\mathrm{Gal}(f/K).$
We know that $\mu$ is defined over $K(X[\ell])$. This allows us to define for each $\sigma\in G_{\ell}$ the isogeny $\sigma(\mu):X\to Y$, which is the Galois-conjugate of $\mu$ (recall that both $X$ and $Y$ are defined over $K$). Then the same construction as in [20, Sect. 4, proof of Prop. 2.4] allows us to define a map
@@
$G_{\ell}=\mathrm{Gal}(K(X[\ell])/K)\twoheadrightarrow\mathrm{Gal}(K(X[1-\delta_{\ell}])/K)=\mathrm{Gal}(f/K)=G$
@@
is a nonzero homomorphism of $\mathrm{Gal}(K)$-modules. Recall that we assume that the Galois module $Y[\ell]$ is trivial. On the other hand, in light of Corollary 3.5, the Galois module $X[\ell]$ admits a filtration, all whose consecutuve quotients are isomorphic to $Q_{\mathfrak{R}_{f}}$. Since $f(x)$ is irreducible over $K$, it follows from Lemma 2.4 that the homomorphism (9) is zero, which is not the case. The obtained contradiction proves that $H$ is not transitive. ∎
@@
So, $n$ is an odd prime, both $f(x)$ and $h(x)\in K[x]$ are degree $n$ polynomials without repeated roots, $f(x)$ is irreducible and $h(x)$ is reducible. Since $n$ is a prime, the reducibility of $h(x)$ implies that the order of $\mathrm{Gal}(h/K)$ is prime to $n$ (see [20, Lemma 2.6]). Let us put $Y=J^{(h,\ell)}$. We are given that the degree $[K(\mathfrak{R}_{h}):K]$ of the field extension $K(\mathfrak{R}_{h})/K$ is not divisible by $n$. Applying Corollary 3.5 to $Y$ and $h(x)$ (instead of $X=J^{(f,\ell)}$ and $f(x)$), we conclude that the degree of the field extension $K(Y[\ell])/K(\mathfrak{R}_{h})$ is either $1$ or a power of $\ell$. This implies that the order of the group $\tilde{G}_{\ell,Y}$ (which is
is prime to $n$, because the prime $\ell$ does not divide $n$. Now the desired result follows readily from Theorem 2.2, because if $r$ is a prime divisor of $n$ then $r=n$.

### v4

@@ 2.1. Galois properties
In this subsection we recall an explicit description of a certain important Galois submodule of $J^{(f,\ell)}[\ell]$ [12, 13] for arbitrary separable $f(x)$, assuming as usual that $\ell$ does not divide $n$.
@@ 3. Isogenous superelliptic jacobians
Let $\ell$ be a prime, $F$ a field of characteristic $\ell$, and $V$ a finite-dimensional vector space over $F$. Let $\mathcal{N}\subset\mathrm{End}_{F}(V)$ be a linear nilpotent Lie subalgebra of $\mathrm{End}_{F}(V)$, and
@@
    $K(X[\ell])/K(\mathfrak{R}_{f})$ is a finite Galois $\ell$-extension, i.e., either $K(X[\ell])=K(\mathfrak{R}_{f})$ or the Galois group $\mathrm{Gal}(K(X[\ell])/K(\mathfrak{R}_{f}))$ is a finite $\ell$-group. In particular, if the order of $\mathrm{Gal}(f/K)$ is prime to $n$ then the order of $\tilde{G}_{\ell,X,K}$ is also prime to $n$.
@@
$\frac{2\mathrm{dim}(J^{(f,\ell)})}{[{\mathbb{Q}}(\zeta_{\ell}):{\mathbb{Q}}]}=\frac{(n-1)(\ell-1)}{(\ell-1)}=n-1.$
@@
is a ring embedding that makes $X[\ell]$ a free ${\mathbb{Z}}[\delta_{\ell}]/\ell$-module of rank $n-1$. This implies that
@@
Now first assertion of (i) follows from Lemma 3.3 applied to
@@
The second one follows readily from the equality
(recall that the prime $\ell$ does not divide $n$).
@@
Lemma 3.7.
Let $\ell$ and $r$ be two distinct odd primes. Let $\mathcal{V}$ be a nonzero finite-dimensional vector space over ${\mathbb{Q}}$. Let $A$ and $B$ be two commuting automorphisms of $\mathcal{V}$ that enjoy the following properties.
  • (i)
    $A^{\ell}=B^{r}=\mathrm{1}_{\mathcal{V}}$ where $\mathrm{1}_{\mathcal{V}}$ is the identity automorphism of $\mathcal{V}$.
  • (ii)
    $A-\mathrm{1}_{\mathcal{V}}$ is an automorphism of $\mathcal{V}$.
  • (iii)
    $B\neq\mathrm{1}_{\mathcal{V}}$.
Then $\mathrm{dim}_{{\mathbb{Q}}}(\mathcal{V})\geq(\ell-1)(r-1)$.
Proof of Theorem 2.2 (modulo Lemmas 3.6 and 3.7).
Recall that $X=J^{(f,\ell)})$.
@@
Both endomorphism algebras $\mathrm{End}^{0}(X)=\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(Y)$ contain an invertible element of multiplicative order $r$. In order to finish the proof of our Theorem, we need to prove the inequality
$\mathrm{dim}_{{\mathbb{Q}}}(\mathrm{End}^{0}(X))\geq(\ell-1)(r-1).$
Let us use Lemma 3.7 applied to $\mathcal{V}=\mathrm{End}^{0}(X)$. We define the automorphisms $A,\ B:\mathcal{V}\to\mathcal{V}$ of the ${\mathbb{Q}}$-vector space $\mathcal{V}=\mathrm{End}^{0}(X)$ as
$v\mapsto\delta_{\ell}v\ \text{ and }v\mapsto vu\quad\forall v\in\mathcal{V}=\mathrm{End}^{0}(X)$
respectively. Clearly, $A$ and $B$ are commuting automorphisms of $\mathcal{V}$ such that both $A^{\ell}$ and $B^{r}$ coincide with the identity automorphism $\mathrm{1}_{\mathcal{V}}$ of $\mathcal{V}$. Since $u$ has multiplicative order $r>1$, $B\neq\mathrm{1}_{\mathcal{V}}$. On the other hand, we know that the ${\mathbb{Q}}$-subalgebra ${\mathbb{Q}}[\delta_{\ell}]$ of $\mathrm{End}^{0}(J^{(f,\ell)})=\mathrm{End}^{0}(X)$ is a subfield with the same identity element as $\mathrm{End}^{0}(X)$. This implies that the nonzero $\delta_{\ell}-\mathrm{1}_{X}$ is an invertible element of the ${\mathbb{Q}}$-algebra $\mathrm{End}^{0}(X)$. It follows that $A-\mathrm{1}_{\mathcal{V}}$ is an invertible automorphism of the ${\mathbb{Q}}$-vector space $\mathcal{V}$. So, $A$ and $B$ satisfy all the conditions of Lemma 3.7. Applying Lemma 3.7, we conclude that $\mathrm{dim}_{{\mathbb{Q}}}(\mathrm{End}^{0}(X))\geq(\ell-1)(r-1)$, which ends the proof of Theorem 2.2.
@@
$G_{\ell}:=\tilde{G}_{\ell,X,K}=\mathrm{Gal}(K(X[\ell])/K),\ G=\mathrm{Gal}(K(X[1-\delta_{\ell}])/K)=\mathrm{Gal}(f/K).$
We know that $\mu$ is defined over $K(X[\ell])$. This allows us to define for each $\sigma\in G_{\ell}$ the isogeny $\sigma(\mu):X\to Y$, which is the Galois-conjugate of $\mu$ (recall that both $X$ and $Y$ are defined over $K$). Then the same construction as in [22, Sect. 4, proof of Prop. 2.4] allows us to define a map
@@
$G_{\ell}=\mathrm{Gal}(K(X[\ell])/K)\twoheadrightarrow\mathrm{Gal}(K(X[1-\delta_{\ell}])/K)=$
$\mathrm{Gal}(K(\mathfrak{R}_{f})/K)=\mathrm{Gal}(f/K)=G$
@@
is a nonzero homomorphism of $\mathrm{Gal}(K)$-modules. Recall that we assume that the Galois action on $Y[\ell]$ is trivial. On the other hand, in light of Corollary 3.5, the Galois module $X[\ell]$ admits a filtration, all whose consecutuve quotients are isomorphic to $Q_{\mathfrak{R}_{f}}$. Since $f(x)$ is irreducible over $K$, it follows from Lemma 2.5 that the homomorphism (9) is zero, which is not the case. The obtained contradiction proves that $H$ is not transitive. ∎
Proof of Lemma 3.7.
Clearly, $B$ is a semisimple (i.e., diagonalizable over $\bar{{\mathbb{Q}}}$) linear operator in $\mathcal{V}$. The same is obviously true for the linear operator $B-\mathrm{1}_{\mathcal{V}}:\mathcal{V}\to\mathcal{V}$. The semisimplicity of $B-\mathrm{1}_{\mathcal{V}}$ implies that $\mathcal{V}$ splits into a direct sum
$\mathcal{V}=\left(B-\mathrm{1}_{\mathcal{V}}\right)(\mathcal{V})\oplus\ker\left(B-\mathrm{1}_{\mathcal{V}}\right)$
of the image $\left(B-\mathrm{1}_{\mathcal{V}}\right)(\mathcal{V})$ and the kernel $\ker\left(B-\mathrm{1}_{\mathcal{V}}\right)$ of $B-\mathrm{1}_{\mathcal{V}}$; clearly, these two subspaces are $B$-invariant. In addition, the restriction of $B-\mathrm{1}_{\mathcal{V}}$ to the subspace
$\mathcal{V}_{0}:=\left(B-\mathrm{1}_{\mathcal{V}}\right)(\mathcal{V})$
is an automorphism of $\mathcal{V}_{0}$. Recall that our conditions on $B$ imply that
$\mathcal{V}_{0}\neq\{0\}.$
Since $A$ and $B$ commute, both subspaces $\left(B-\mathrm{1}_{\mathcal{V}}\right)(\mathcal{V})=\mathcal{V}_{0}$ and $\ker\left(B-\mathrm{1}_{\mathcal{V}}\right)$. are also $A$-invariant. Let
$A_{0},\ B_{0}:\mathcal{V}_{0}\to\mathcal{V}_{0}$
be the restrictions to $\mathcal{V}_{0}$ of $A$ and $B$ respectively. Clearly, $A_{0}$ and $B_{0}$ commute, and both $A_{0}^{\ell}$ and $B_{0}^{r}$ coincide with the identity automorphism $\mathrm{1}_{\mathcal{V}_{0}}$ of $\mathcal{V}_{0}$. In addition, both $A_{0}-\mathrm{1}_{\mathcal{V}_{0}}$ and $B-\mathrm{1}_{\mathcal{V}_{0}}$ are automorphisms of $\mathcal{V}_{0}$.
Let
$\mathcal{P}_{A_{0}}(t),\ \mathcal{P}_{B_{0}}(t)\in{\mathbb{Q}}[t]$
be the minimal polynomials of $A_{0}:\mathcal{V}_{0}\to\mathcal{V}_{0}$ and $B_{0}:\mathcal{V}_{0}\to\mathcal{V}_{0}$ respectively. Both minimal polynomials are monic of positive degree, and all their coefficients are rational numbers. Clearly, $\mathcal{P}_{A_{0}}(t)$ divides $t^{\ell}-1$ and $\mathcal{P}_{B_{0}}(t)$ divides $t^{r}-1$; in addition, $t-1$ divides neither $\mathcal{P}_{A_{0}}(t)$ nor $\mathcal{P}_{B_{0}}(t)$. Recall that
$t^{\ell}-1=(t-1)\Phi_{\ell}(t),\quad t^{r}-1=(t-1)\Phi_{\ell}(t)$
where $\Phi_{\ell}(t)$ and $\Phi_{r}(t)$ are $\ell$th and $r$th cyclotomic polynomials respectively; they both are irreducible over ${\mathbb{Q}}$. It follows that
$\mathcal{P}_{A_{0}}(t)=\Phi_{\ell}(t),\quad\mathcal{P}_{B_{0}}(t)=\Phi_{r}(t).$
This implies that the ${\mathbb{Q}}$-subalgebra ${\mathbb{Q}}[A_{0}]$ of $\mathrm{End}_{{\mathbb{Q}}}(\mathcal{V}_{0})$ generated by $A_{0}$ is isomorphic to the $\ell$th cyclotomic field
${\mathbb{Q}}[t]/\Phi_{\ell}(t){\mathbb{Q}}[t]\cong{\mathbb{Q}}(\zeta_{\ell}).$
Similarly, the ${\mathbb{Q}}$-subalgebra ${\mathbb{Q}}[B_{0}]$ of $\mathrm{End}_{{\mathbb{Q}}}(\mathcal{V}_{0})$ generated by $B_{0}$ is isomorphic to the $r$th cyclotomic field
${\mathbb{Q}}[t]/\Phi_{r}(t){\mathbb{Q}}[t]\cong{\mathbb{Q}}(\zeta_{r}).$
Since $A_{0}$ and $B_{0}$ commute, the (commutative) ${\mathbb{Q}}$-subalgebras ${\mathbb{Q}}[A_{0}]$ and ${\mathbb{Q}}[B_{0}]$ also commute. This implies that the nonzero ${\mathbb{Q}}$-vector space $\mathcal{V}_{0}$ carries the natural structure of a module over the ${\mathbb{Q}}$-algebra
${\mathbb{Q}}[A_{0}]\otimes_{{\mathbb{Q}}}{\mathbb{Q}}[B_{0}]\cong{\mathbb{Q}}(\zeta_{\ell})\otimes_{{\mathbb{Q}}}{\mathbb{Q}}(\zeta_{r}).$
Since $r$ and $\ell$ are distinct odd primes, the cyclotomic fields ${\mathbb{Q}}(\zeta_{\ell})$ and ${{\mathbb{Q}}}{\mathbb{Q}}(\zeta_{r})$ are linearly disjoint over ${\mathbb{Q}}$; actually, this tensor product is canonically isomorphic to $\ell r$th cyclotomic field ${\mathbb{Q}}(\zeta_{\ell r})$ of degree $(\ell-1)(r-1)$. It follows that $\mathcal{V}_{0}$ carries the natural structure of a ${\mathbb{Q}}(\zeta_{\ell r})$-vector space. Hence, $\mathrm{dim}_{{\mathbb{Q}}}(\mathcal{V}_{0})$ is divisible by the degree
$[{\mathbb{Q}}(\zeta_{\ell r}):{\mathbb{Q}}]=(\ell-1)(r-1).$
Since $\mathcal{V}_{0}\neq\{0\}$, we conclude that $\mathrm{dim}_{{\mathbb{Q}}}(\mathcal{V}_{0})\geq(\ell-1)(r-1).$ Taking into account that $\mathcal{V}_{0}$ is a subspace of $\mathcal{V}$, we conclude that
$\mathrm{dim}_{{\mathbb{Q}}}(\mathcal{V})\geq\mathrm{dim}_{{\mathbb{Q}}}(\mathcal{V}_{0})\geq(\ell-1)(r-1).$
This ends the proof of our Lemma. ∎
@@
So, $n$ is an odd prime, both $f(x)$ and $h(x)\in K[x]$ are degree $n$ polynomials without repeated roots, $f(x)$ is irreducible and $h(x)$ is reducible. Since $n$ is a prime, the reducibility of $h(x)$ implies that the order of $\mathrm{Gal}(h/K)$ is prime to $n$ (see [22, Lemma 2.6]). Let us put $Y=J^{(h,\ell)}$. We are given that the degree $[K(\mathfrak{R}_{h}):K]$ of the field extension $K(\mathfrak{R}_{h})/K$ is not divisible by $n$. Applying Corollary 3.5 to $Y$ and $h(x)$ (instead of $X=J^{(f,\ell)}$ and $f(x)$), we conclude that the order of the group $\tilde{G}_{\ell,Y,K}$ is prime to $n$. Now the desired result follows readily from Theorem 2.2, because if $r$ is a prime divisor of $n$ then $r=n$.
5. Doubly transitive and cyclic Galois groups
Proof of Corollary 1.4.
By definition of the field
$K_{h}:=K(\mathfrak{R}_{h}),$
the polynomial $h(x)$ splits into a product of linear factors over $K_{h}$. Recall that $\mathrm{Gal}(K_{h}/K)=\mathrm{Gal}(h/K)$ is a cyclic group of prime order $n$ and $\mathrm{Gal}(f/K)=\mathrm{Gal}(K(\mathfrak{R}_{f})/K)$ is doubly transitive. In light of Proposition 1.8 of [22], the field extensions $K(\mathfrak{R}_{f})/K$ and $K_{h}/K$ are linearly disjoint. This implies that $f(x)$ remains irreducible over $K_{h}$. Now the desired result follows readily from Theorem 1.1 applied to $f(x)$ and $h(x)$ over $K_{h}$ (instead of $K$). ∎